OpenStudy (anonymous):

Helpppppppp. This is hard to me if x^2 + x - 3 = 0 what is the value of x^3 + 4x^2 + 2014 ?

jimthompson5910 (jim_thompson5910):

First I recommend solving x^2 + x - 3 = 0 for x You would use the quadratic formula to solve for x

jimthompson5910 (jim_thompson5910):

Once you have your solutions, plug them into x^3 + 4x^2 + 2014 one at a time

OpenStudy (anonymous):

is there an alternative way without solve for x ?

OpenStudy (anonymous):

i make that equation be x^2 + x - 3 = 0 times by x on both sides, i get x^3 + x^2 - 3x = 0 but i dont know how to make this be like equation x^3 + 4x^2 + 2014 ?

OpenStudy (anonymous):